Let l be any positive integer. Use sandwich theorem for sequences to prove that the sequence {1/n^ l} converges to zero.


1
Expert's answer
2021-01-26T04:29:31-0500

We consider the zero sequence, that is the sequence "(0)_{n=1}^{n=\infty}"

We have that for a positive integer "l, 0 \leq 1/n^l" ".... *"

Also, we have that "n \leq n^l"

"\therefore 1/n^l \leq 1/n" ".... **"

Combining * and **, we have "0 \leq 1/n^l \leq 1/n"

Since "\lim_{n \to \infty }0 = 0 = \lim_{n \to \infty} 1/n"

By sandwich theorem, we have that

"\lim_{n \to \infty} 1/n^l = 0" as desired.